OpenStudy (anonymous):

factor: 4x^2-21x+5=0 please/thanks!

OpenStudy (espex):

It would be easiest to use the quadratic formula for this.

OpenStudy (rulnick):

(4x-1)(x-5)=0.

OpenStudy (rulnick):

So if you need to solve, the two solutions would come from 4x-1 = 0 and x-5 = 0.

OpenStudy (anonymous):

Thank You!

OpenStudy (rulnick):

welcome